Junaid was 28 when his daughter Hafsa was born. If junaid is now three times as old as hafsa,how many years ago was he 5 times as old as Hafsa?
Answer by ankor@dixie-net.com(22740)

Junaid was 28 when his daughter Hafsa was born.
If junaid is now three times as old as hafsa, how many years ago was he 5 times as old as Hafsa?
:
From the information given we know Jun is 28 yrs older than Haf, therefore
j = h + 28
:
Jun is now 3 times as old as h, therefore
j = 3h
replace j with (h+28)
h + 28 = 3h
28 = 3h - h
28 = 2h
h = 28/2
h = 14 yrs is Haf's present age
then
14 + 28 = 42 yrs is Jun's present age
:
"how many years ago was he 5 times as old as Hafsa?"
let y = no. of yrs ago that this is true
42 - y = 5(14-y)
42 - y = 70 - 5y
-y + 5y = 70 - 42
4y = 28
y = 28/4
y = 7 yrs ago Jun was 5 time Haf's age
:
:
See if this checks out
7yrs ago:
j = 35
h = 7; confirms our solutions